OpenStudy (anonymous):

Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

\[\sqrt[3]{5}*\sqrt{5} \over \sqrt[3]{5^5}\]

OpenStudy (anonymous):

0.2615320972

OpenStudy (anonymous):

can you turn that into a 5 with a fraction?

OpenStudy (anonymous):

like 5 then the fracton?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

need anything else

OpenStudy (anonymous):

yes actually i'm sorry

OpenStudy (anonymous):

tellme

OpenStudy (anonymous):

More math

OpenStudy (anonymous):

Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

\[(1,8) and (4,512)\]

OpenStudy (anonymous):

Sorry wrong question!